Need help on the second part of this word problem. Showing steps would be helpful...Thank you :)
Construction of the Tower of Pisa was completed in 1360. By 1990, the tilt of the tower was so severe that it was closed for renovation. Renovators were able to reduce the tower's 1990 tilt by 17 inches. The resultant tower leans 13.5 feet (162 inches) off the perpendicular. When the tower was reopened in 2001, officials forecast that it would take 300 years for the tower to return to its 1990 tilt.†
(a) Construct a linear formula that models the lean of the renovated tower, where l is the number of inches from the perpendicular and t is the number of years since 2001.
The formula is: 162+(17/300)t

Use the formula from part (a) to predict the lean of the tower in 2150. (Round your answer to two decimal places.)

Answer by rothauserc(4718)

l = 162+(17/300)t
t = 2150 - 2001 = 149, then
l = 162 + (17/300)149
l = 162 + 8.443333333
l = 170.443333333
therefore,
l = 170.44 inches